A model for the path of a toy rocket is given by 68t-4.9t(squared), where h is the altitude in meters and t is the time in seconds. Explain how to find both the maximum altitude of the rocket and how long it takes to reach that altitude.
Answer by Alan3354(69443)

It's a parabola. h(t) = -4.9t^2 + 68t
The time to max height is the line of symmetry, -b/2a
t = -b/2a = -68/-9.8
t = 340/49 seconds = ~ 6.939 seconds
-----------
Max ht = h(340/49) = 566440/2401 meters
=~ 235.9 meters
------------------